OpenStudy (anonymous):

Find the constant necessary to make a perfect square trinomial. 2c^2+7c+__

OpenStudy (anonymous):

49/4 right

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

if you want to write it as a perfect square you have to factor out the 2 first and start with \[2(x^2+\frac{7}{2}x)\]

OpenStudy (anonymous):

take my medals

OpenStudy (anonymous):

and then square it

OpenStudy (misty1212):

then half of \(\frac{7}{2}\)is \(\frac{7}{4}\) square that and get \[\frac{49}{16}\]

OpenStudy (misty1212):

the square will be \[2(x+\frac{7}{4})^2\]

OpenStudy (anonymous):

you are correct